Danny can finish a job in 10 days. Together, Miguel and Danny can finish the same job in 6 days. How many days will it take Miguel to finish the job alone?
Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

Let x be the number of days Miguel takes to finish the job alone, then
:
1/10) + (1/x) = 1/6
:
Note that the above equation represents the contribution of Miguel and Danny working together
:
The lowest common denominator is 30x
:
(3x + 30) / 30x = 5x / 30x
:
multiply both sides by 30x
:
3x + 30 = 5x
:
2x = 30
:
x = 15
:
*************************************************
it takes Miguel 15 days to complete the job alone
*************************************************
:

Answer by ikleyn(52858)

.
For a wide variety of similar solved joint-work problems with detailed explanations see the lessons
    - Using Fractions to solve word problems on joint work,
    - Solving more complicated word problems on joint work,
    - Selected joint-work word problems from the archive
in this site.

Read them and get be trained in solving joint-work problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Rate of work and joint work problems" of the section "Word problems".